Quiz 3

\underline{Problem 1}: Consider a distillation column with a feed of 10 gmol/s benzene, 20 gmol/s o-xylene and 15 gmol/s toluene at 101.33 kPa and 383 K.

• For a benzene recovery of 98%, what is the minimum number of trays if the ratio of benzene to o-xylene in the overhead is 100? Find the bottoms and tops compositions.

• Periodically a small amount of H_{2}S appears in the feed. Since it is undesirable to have this component in the product, explain qualitatively how you would design and operate this column to separate the H_{2}S.

• An overhead product of 55% benzene, 40% toluene and 5% o-xylene is recovered as saturated liquid. Can cooling water at 310 K be used if the column operates at 200 kPa?

Antione Equation:

log_{10}(P) = A − \frac{B}{T + C}

where P is the vapor pressure in kPa and T is the temperature in K.

Compound A  B  C
Benzene  6.06832 1236.034 -48.990
Toluene  6.12072 1374.901 -15.440
o-Xylene  6.13132 1480.155 -58.804